What is the angle between the following two vectors: [111] and [010] Ο 60° Ο 30° Ο 54.70 Ο 45ο
Answer #1

Ans :

  \small \dot{u} = [ 1 1 1 } &  \small \dot{v} = [ 0 1 0]

Angle between two vectors \small \ddot{} \small \dot{u} & \small \dot{v} is given by

cos \small \Theta = u .v / II\small \dot{u}II II \small \dot{v} II

II\small \dot{u}II = Magnitude of  \small \dot{u} = \small \sqrt{1^{2}+1^{2}+1^{2}} =  \small \sqrt{3}  

II\small \dot{v}II = Magnitude of \small \dot{v}    = \small \sqrt{0^{2}+1^{2}+0^{2}} = \small \sqrt{1} = 1

Dot product of  \small \dot{u} & \small \dot{v}

u .v = (1 x1x1 ) + (0 x 1 x0)  

= 1 + 0 = 1

Using formula for angle

cos \small \Theta = 1 / ( \small \sqrt{3} x  1 )

= 1 / \small \sqrt{3}

Therefore \small \Theta = cos-1  (1 / \small \sqrt{3} )

= 54.7350

Angle between two vectors = 54.70  ( Rounding off )